# Amortized Bayesian Meta-Learning for Low-Rank Adaptation of Large Language Models

### v2

@@ 4 Method
$\hskip 0.0ptp(\theta)$ $=p(\mu_{\theta},\sigma_{\theta})=\mathcal{N}(\mu_{\theta};0,\mathbf{I})\cdot\text{Gamma}\left(\frac{1}{\sigma_{\theta}^{2}};a_{0},b_{0}\right),$
$\hskip 0.0pt\text{KL}(q(\theta)\,\|\,p(\theta))$ $=-\log p(\theta).$
@@
  Run a step gradient descent to minimize w.r.t $\phi_{i}^{c}$$\mathbb{E}_{q_{\theta^{c}}(\phi_{i}^{c}|D_{i})}[\log p(D_{i}|\phi_{i}^{c})]+\beta\text{KL}\big(q_{\theta^{c}}(\phi_{i}^{c}|D_{i})\big|\big|p(\phi_{i}^{c}|\theta^{c})\big).$
@@
Lastly, $p(D_{i}|\phi_{i})$ is defined as the joint probability assigned to $D_{i}$ where the LLM takes $\phi_{i}$ as its weights. The trainable parameters are the LoRA adapters $\mathbf{A}$ and $\mathbf{B}$. However, we introduce four pairs of these adapters to compute both the mean and variance of the LoRA outputs on local and global model weights. $\mathbf{I}$ is identity matrix, and $c$ is a hyperparameter constant dependent on the spread of pretrained LLM weights. $a_{0}$ and $b_{0}$ are hyperparameters, and the simplification of the KL term as $\log p(\theta)$ follows ravi2018amortized.
Balancing the reconstruction error.
LLMs are often overparameterized. As a result, probabilistic quantities on the space of weights, $\text{KL}\big(q_{\theta}(\phi_{i}|D_{i})\big|\big|p(\phi_{i}|\theta)\big)$ and $\text{KL}(q(\theta)||p(\theta))$, can overwhelm quantities on the data space, $\log p(D_{i}|\phi_{i})$. $\beta$-VAE (Higgins2016betaVAELB) and Bayesian neural network approaches (trinh2022tackling) introduce hyperparameters to temper the likelihood versus regularization terms. Inspired by this idea, we introduce hyperparameter $\beta$, resulting in the following objective,
5 Empirical Evaluations
5.1 Few-Shot Learning
We first examine models fine-tuned by our ABMLL approach as few-shot learners on unseen tasks in natural text datasets.
5.1.1 Experimental Setup
Model and datasets.
We fine-tune Llama3-8B on CrossFit (ye-etal-2021-crossfit) and UnifiedQA (ye-etal-2021-crossfit), two text datasets commonly used to train meta-learning models.
We train and evaluate in three settings. First, we use cls-45 (ye-etal-2021-crossfit), where models are trained on classification tasks, and evaluated on other distinct classification tasks. Second, we use cls-23 (ye-etal-2021-crossfit), where models are evaluated on the same classification tasks as in cls-45, but are trained on a mix of classification and other tasks including question-answering and natural language inference. Finally, following min-etal-2022-metaicl, we test stronger generalizations on more narrowly defined tasks, where models train on other tasks, but evaluate on only natural language inference (NLI), paraphrasing (Para), and knowledge-based multiple choice questions-answers (MCQA), respectively.
Metrics.
Baselines.
We use four baseline methods that can viably scale to Llama3-8B. Pretrained is the off-the-shelf LLM. Regular LoRA is the default LoRA method trained on the whole randomly shuffled training dataset. Structured LoRA also uses the default LoRA, but the training dataset follows the same “structure” as our method: it is iteratively trained 5 gradient steps on one task at a time, dropping the global variable (and dropping the reinitialization of the task-specific variable to the global variable). Thus, it tests the effect of our generative model on performance. Reptile (Nichol2018OnFM) implements the Reptile meta-learning algorithm.
Implementation details.
All experiments run 500 epochs with a single A100 GPU with 40GB of memory. All methods use the AdamW optimizer (loshchilov2017decoupled), batch-size of 2, inner-loops with 5 gradient steps, LoRA adapters with rank =8=8 following the standard practice, and learning rate is tuned in $10^{-6},10^{-4}]$}. For ABMLL, $\beta=10^{-8}$}, $c=e^{-20}$}. For the gamma prior, $a_{0}=1,b_{0}=0.01$} following ravi2018amortized. It takes one sample from the reparameterization step during inference. During validation on the unseen dataset, all models train 5 gradient steps on 5 batches from this dataset and evaluate on the rest.
5.1.2 Experimental Results
Table 1 reports test scores for each model from three random seeds, where the test results are taken from each model’s best validation accuracy epoch. Statistically significant best performances, including ties, are in bold. On cls-45 and cls-23, ABMLL performs best on both metrics, suggesting that ABMLL trains a general learner able to adapt to a variety of problems. The performance of ABMLL is tied with Reptile on tasks with more distinct train-evaluation differences, possibly because the tested ability is more difficult to acquire through training tasks, bringing the performance of these two meta-learning models closer together. Meanwhile, the two meta-learning methods perform significantly better than the rest, suggesting that meta-learning is an important ingredient for fine-tuning LLMs with stronger generalization.
5.1.3 Memory Consumption
An advantage of ABMLL over many meta-learning methods is scalability. Although ABMLL introduces four pairs of LoRA adapters, pretrained weights from both ABMLL and LoRA still need to be computed during a forward pass, despite having no gradients attached to them. We compute peak memory during fine-tuning on the cls-45 dataset and find that ABMLL only requires $7.6\%$ more memory than regular LoRA (25.6 GB for ABMLL compared to 23.8 GB for regular LoRA).
Table 1: Test accuracy and ECE across three random seeds, with standard error. Statistically significant best performances, including tied ones, are bolded.
Method cls-45 Acc $\uparrow$ cls-23 Acc $\uparrow$ NLI Acc $\uparrow$ Para Acc $\uparrow$ MCQA Acc $\uparrow$
Pretrained $26.1\%$ $\pm 0.1\%$ $26.0\%$ $\pm 0.1\%$ $57.6\%$ $\pm 0.0\%$ $57.0\%$ $\pm 0.0\%$ $71.8\%$ $\pm 0.1\%$
Method cls-45 ECE $\downarrow$ cls-23 ECE $\downarrow$ NLI ECE $\downarrow$ Para ECE $\downarrow$ MCQA ECE $\downarrow$
Pretrained $0.458$ $\pm 0.000$ $0.458$ $\pm 0.000$ $0.419$ $\pm 0.000$ $0.430$ $\pm 0.000$ $0.276$ $\pm 0.000$
5.2 Model Pruning
Model pruning is a way to improve the efficiency of LLMs by reducing their size and computational requirements (Ashkboos2024SliceGPTCL; Sun2023ASA). It is also a measurement of model robustness as it tests model performance by removing redundant parameters, potentially dropping spurious correlations. Bayesian neural networks are known to be resource-efficient (blundell_2015). Inspired by this idea, we test the performance of ABMLL and benchmarks by setting a percentage of neurons in each layer embedding to zero, sorted by magnitudes of these neurons.
Table 3(b) shows that ABMLL is significantly more robust against this form of pruning than the other methods. This result suggests that ABMLL is robust and reliable, learning generalizable features that are not as tied to specific parameters. Although regularization methods for other benchmarks such as weight-decay can potentially improve performance under pruning, we make the following observations: (1) weight-decay in general worsens performance under the no-pruning scenario, and (2) weight-decay achieves better performance than no-regularization under pruning but is still worse than ABMLL.
Method $0\%$ Pruned $1\%$ Pruned $10\%$ Pruned $20\%$ Pruned $30\%$ Pruned
(a) NLI.
Method $0\%$ Pruned $1\%$ Pruned $10\%$ Pruned $20\%$ Pruned $30\%$ Pruned
(b) Para.
5.3 Ablation Studies
In this section, we analyze the effect of different terms in our training objective, Equation 2.
Table 3 shows performance of ABMLL with different values of $\beta$ on cls-45. Results show that $\log_{10}\beta=0$, i.e., $\beta=1$, drops performance significantly. Thus, it is critical to balance reconstruction error and the KL terms that control how close the task-specific parameters $\phi_{i}$ are to global parameters $\theta$, and how close $\theta$ are to the prior $p(\theta)$.
At the other extreme where the KL terms are too tempered, the objective becomes degenerate, leading to poor performance ($\log_{10}\beta=-16$). In our experiments we searched for $\log_{10}\beta\in\{0,4,5,6,7,8,9,10\}$ with one random seed on cls-45 and identified the optimal value as $\log_{10}\beta=-8$. We use this setting for $\beta$ across all experiments, including Table 1.
Table 3: Validation accuracy and ECE across values of $\beta$ on cls-45. Lower values of $\beta$ means that the KL terms are more tempered, leading to pure maximization of data log likelihood on the objective in the limit.
$\log_{10}\beta$ Accuracy $\uparrow$ ECE $\downarrow$
0 $64.5\%$ $0.395$
-5 $70.4\%$ $0.289$
-8 $75.2\%$ $0.262$
-16 $60.6\%$ $0.395$
6 Discussion
Our results show that Bayesian modeling with meta-learning for LoRA fine-tuning achieves strong performance, measured in both accuracy and robustness. Additionally, ablation studies demonstrate the importance of the $\beta$ hyperparameter to achieving proper balancing of the training objective. These studies also verify the validity of this objective, since tempering the KL terms too heavily leads to poor performance. In the remainder of the paper we consider the implications of these results for broader questions about Bayesian methods and inductive biases for large language models. We also identify some of the limitations of our analyses and directions for future work.
Bridging the gap between Bayesian deep learning and LLMs.
Bayesian probabilistic modeling provides a principled way to incorporate human knowledge into models or to quantify uncertainty (Blei2014BuildCC; Griffiths2008). In the form of Bayesian neural networks, it has been elegantly explored for smaller neural networks where uncertainty is added to their weights (MACKAY199573; blundell_2015). The era of large models poses a challenge in connecting Bayesian methods with deep learning because of their large computational requirements. Although Bayesian methods have the potential to improve LLM interpretability, uncertainty quantification, and adaptation to new domains, they are currently under-explored (pmlr-v235-papamarkou24b).
Our work bridges this gap by using LoRA to make LLMs amenable to Bayesian probabilistic modeling. Additionally, we use the paradigm of meta-learning to develop a generative probabilistic model that offers a novel and effective way to conduct LLM fine-tuning.
Inductive bias in the era of LLMs.
Interpreting LLMs is challenging; incorporating inductive bias into their training or fine-tuning is even more so. Meta-learning is a way to incorporate inductive biases into LLMs, from gradient-based approaches (mamlenllm; lift2025) to in-context learning (min-etal-2022-metaicl; chen-etal-2022-meta). As a different approach, mccoy_inductive constructs synthetic datasets and uses meta-learning to enforce grammatical awareness in language models prior to training on natural corpora. However, all of these methods are difficult to scale to larger models that have at least several billion parameters. Our work provides another approach where amortized Bayesian inference captures a favorable inductive bias by modeling task-specific variables as being generated from global variables. Critically, this approach has memory overhead that is constant with respect to the number of tasks.
Limitations and future work.
We view ABMLL as filling crucial gaps between Bayesian methods and LLMs. While its performance is comparable to the other scalable meta-learning method, Reptile, our results show that ABMLL has greater robustness under model pruning, similar to other findings in Bayesian neural networks.
There are several directions for future work. It would be valuable to test these meta-learning methods on more models and compare their effectiveness versus model sizes. Additionally, an advantage of Bayesian methods, as well as inductive bias distillation, is that they may require less data. It would be interesting to adopt ABMLL in a limited data regime and test its performance.
Conclusion.
Meta-learning is an effective method for supporting better generalization across datasets, but its demands on computation and memory can make it difficult to apply to large language models. We have shown how meta-learning can be used to adapt LLMs by combining Amortized Bayesian Meta-Learning with Low-Rank Adaptation. This approach results in greater accuracy and across several benchmarks, and produces models that are more robust to pruning than existing methods for meta-learning in LLMs.
Acknowledgments
We thank Qinyuan Ye for helpful discussion. This work was supported by grant N00014-23-1-2510 from the Office of Naval Research.
Appendix A Appendix
A.1 Additional Experiment Results

### v3

@@ 4 Method
$\begin{array}[]{cc}\begin{aligned} \mu_{\theta}&=\mathbf{B}_{\mu_{\theta}}\mathbf{A}_{\mu_{\theta}},\\ \log\sigma_{\theta}^{2}&=\mathbf{B}_{\sigma_{\theta}}\mathbf{A}_{\sigma_{\theta}}+c\mathbf{I},\\ \mu_{\phi}&=\mathbf{B}_{\mu_{\phi}}\mathbf{A}_{\mu_{\phi}},\\ \log\sigma_{\phi}^{2}&=\mathbf{B}_{\sigma_{\phi}}\mathbf{A}_{\sigma_{\phi}}+c\mathbf{I}\end{aligned}&\begin{aligned} p(\phi_{i}\mid\theta)&=\mathcal{N}(\phi_{i};\mu_{\theta}+\mathbf{W_{0}},\sigma_{\theta}^{2}),\\ q_{\theta}(\phi_{i}\mid D_{i})&=\mathcal{N}(\phi_{i};\mu_{\phi}+\mathbf{W_{0}},\sigma_{\phi}^{2}),\\ p(\theta)&=p(\mu_{\theta},\sigma_{\theta})=\mathcal{N}(\mu_{\theta};0,\mathbf{I})\cdot\text{Gamma}\left(\frac{1}{\sigma_{\theta}^{2}};a_{0},b_{0}\right),\\ \text{KL}(q(\theta)&\,\|\,p(\theta))=-\log p(\theta)\end{aligned}\end{array}$
Lastly, $p(D_{i}|\phi_{i})$ is defined as the joint probability assigned to $D_{i}$ where the LLM takes $\phi_{i}$ as its weights. The trainable parameters are the LoRA adapters $\mathbf{A}$ and $\mathbf{B}$. However, we introduce four pairs of these adapters to compute both the mean and variance of the LoRA outputs on local and global model weights. $\mathbf{I}$ is identity matrix, and $c$ is a hyperparameter constant dependent on the spread of pretrained LLM weights. $a_{0}$ and $b_{0}$ are hyperparameters, and the simplification of the KL term as $\log p(\theta)$ follows Ravi and Beatson (2019).
Balancing the reconstruction error.
LLMs are often overparameterized. As a result, probabilistic quantities on the space of weights, $\text{KL}\big(q_{\theta}(\phi_{i}|D_{i})\big|\big|p(\phi_{i}|\theta)\big)$ and $\text{KL}(q(\theta)||p(\theta))$, can overwhelm quantities on the data space, $\log p(D_{i}|\phi_{i})$. $\beta$-VAE (Higgins et al., 2016) and Bayesian neural network approaches (Trinh et al., 2022) introduce hyperparameters to temper the likelihood versus regularization terms. Inspired by this idea, we introduce hyperparameter $\beta$, resulting in the following objective,
5 Empirical Evaluations
5.1 Performance on Few-Shot Learning Datasets
We examine few-shot learning performance by meta-learning methods on natural corpora.
5.1.1 Experimental Setup
Model and datasets.
We fine-tune Llama3-8B and Qwen2-7B on CrossFit and UnifiedQA (Ye et al., 2021), two text datasets commonly used to train meta-learning models.
We train and evaluate in three settings. The first setting is cls-45, where models are trained on classification tasks, and evaluated on other distinct classification tasks. The second setting is cls-23, where models are evaluated on the same classification tasks as in cls-45, but are trained on a mix of classification and other tasks including question-answering and natural language inference. Finally, following Min et al. (2022), we test stronger generalizations on more narrowly defined tasks, where models train on other tasks, but evaluate on only natural language inference (NLI), paraphrasing (Para), and knowledge-based multiple choice questions-answers (MCQA), respectively.
Metrics.
Baselines.
We use five baseline methods that can viably scale to our context. Pretrained is the off-the-shelf LLM. Regular LoRA is the default LoRA method trained on the whole randomly shuffled training dataset. Structured LoRA also uses the default LoRA, but the training dataset follows the same “structure” as our method: it is iteratively trained 5 gradient steps on one task at a time, dropping the global variable. Thus, it tests the effect of our generative model on performance. Reptile (Nichol et al., 2018) implements the Reptile meta-learning algorithm.
Implementation details.
ABMLL takes one sample from the reparameterization step during inference. During validation on the unseen dataset, all models perform 5-shot learning: training 5 gradient steps on 5 batches from this dataset and evaluating on the rest. Appendix A.3 includes more details.
5.1.2 Experimental Results - CrossFit and UnifiedQA
Meta-learning outperforms its counterparts without meta-learning.
Table 1 reports test scores for each model from three random seeds, where the test results are taken from each model’s best validation accuracy epoch. For methods without in-context examples, ABMLL performs best on both metrics on cls-45 and cls-23, suggesting that ABMLL trains a general learner able to adapt to a variety of problems. The performance of ABMLL is tied with Reptile on tasks with more distinct train-evaluation differences, possibly because the tested ability is more difficult to acquire through training tasks, bringing the performance of these two meta-learning models closer together. Meanwhile, the two meta-learning methods perform significantly better than the rest, suggesting that meta-learning is an important ingredient for fine-tuning LLMs with stronger generalization.
Memory consumption.
An advantage of ABMLL over many meta-learning methods is scalability. Although ABMLL introduces four pairs of LoRA adapters, pretrained weights from both ABMLL and LoRA still need to be computed during a forward pass, despite having no gradients attached to them. ABMLL increases peak memory by only $7.6\%$ over regular LoRA on cls-45.
Table 1: Test accuracy and ECE across three random seeds, with standard error, for two models (Llama3 and Qwen2). Best performance is shown in bold.
Model Method cls-45 Acc $\uparrow$ cls-23 Acc $\uparrow$ NLI Acc $\uparrow$ Para Acc $\uparrow$ MCQA Acc $\uparrow$
Llama3 Pretrained $26.1\%$ $\pm 0.1\%$ $26.0\%$ $\pm 0.1\%$ $57.6\%$ $\pm 0.0\%$ $57.0\%$ $\pm 0.0\%$ $71.8\%$ $\pm 0.1\%$
Qwen2 Pretrained $60.5\%$ $\pm 0.1\%$ $60.5\%$ $\pm 0.1\%$ $69.3\%$ $\pm 0.1\%$ $57.1\%$ $\pm 0.1\%$ $77.5\%$ $\pm 0.2\%$
Regular LoRA $66.3\%$ $\pm 0.3\%$ $66.0\%$ $\pm 0.2\%$ $79.7\%$ $\pm 0.3\%$ $63.3\%$ $\pm 0.9\%$ $77.5\%$ $\pm 0.2\%$
Struct. LoRA $69.6\%$ $\pm 0.2\%$ $67.4\%$ $\pm 0.2\%$ $79.1\%$ $\pm 0.2\%$ $62.2\%$ $\pm 0.4\%$ $77.8\%$ $\pm 0.2\%$
Reptile $72.9\%$ $\pm 0.1\%$ $70.0\%$ $\pm 0.4\%$ $82.6\%$ $\pm 0.5\%$ $63.4\%$ $\pm 0.2\%$ $79.0\%$ $\pm 0.1\%$
ABMLL (ours) $74.1\%$ $\pm 0.7\%$ $72.4\%$ $\pm 0.3\%$ $83.6\%$ $\pm 0.0\%$ $66.1\%$ $\pm 0.9\%$ $77.2\%$ $\pm 0.7\%$
Model Method cls-45 ECE $\downarrow$ cls-23 ECE $\downarrow$ NLI ECE $\downarrow$ Para ECE $\downarrow$ MCQA ECE $\downarrow$
Llama3 Pretrained $0.458$ $\pm 0.000$ $0.458$ $\pm 0.000$ $0.419$ $\pm 0.000$ $0.430$ $\pm 0.000$ $0.276$ $\pm 0.000$
Qwen2 Pretrained $0.397$ $\pm 0.000$ $0.397$ $\pm 0.000$ $0.298$ $\pm 0.000$ $0.428$ $\pm 0.000$ $0.223$ $\pm 0.000$
Regular LoRA $0.322$ $\pm 0.001$ $0.330$ $\pm 0.001$ $0.269$ $\pm 0.001$ $0.423$ $\pm 0.001$ $0.241$ $\pm 0.002$
Struct. LoRA $0.307$ $\pm 0.000$ $0.321$ $\pm 0.001$ $0.270$ $\pm 0.002$ $0.401$ $\pm 0.002$ $0.239$ $\pm 0.002$
Reptile $0.279$ $\pm 0.001$ $0.293$ $\pm 0.002$ $0.231$ $\pm 0.003$ $0.373$ $\pm 0.001$ $0.233$ $\pm 0.002$
ABMLL (ours) $0.273$ $\pm 0.009$ $0.285$ $\pm 0.006$ $0.208$ $\pm 0.008$ $0.364$ $\pm 0.004$ $0.249$ $\pm 0.004$
Table 2: Test accuracy and ECE across three random seeds, with standard error, for two models (Llama3 and Qwen2) in the ICL setting. Best performances is shown in bold.
Model Method cls-45 Acc $\uparrow$ cls-23 Acc $\uparrow$ NLI Acc $\uparrow$ Para Acc $\uparrow$ MCQA Acc $\uparrow$
Llama3 MetaICL $71.0\%$ $\pm 0.1\%$ $70.8\%$ $\pm 0.2\%$ $83.3\%$ $\pm 0.3\%$ $61.6\%$ $\pm 0.2\%$ $80.1\%$ $\pm 0.2\%$
Reptile-MetaICL $74.2\%$ $\pm 0.8\%$ $73.5\%$ $\pm 0.2\%$ $83.6\%$ $\pm 0.2\%$ $64.0\%$ $\pm 0.8\%$ $80.2\%$ $\pm 0.2\%$
ABMLL-MetaICL (ours) $77.4\%$ $\pm 0.5\%$ $75.4\%$ $\pm 0.6\%$ $85.2\%$ $\pm 0.1\%$ $68.1\%$ $\pm 0.3\%$ $80.2\%$ $\pm 0.3\%$
Qwen2 MetaICL $72.2\%$ $\pm 0.5\%$ $70.2\%$ $\pm 0.3\%$ $83.8\%$ $\pm 0.2\%$ $76.2\%$ $\pm 0.2\%$ $81.3\%$ $\pm 0.1\%$
Reptile-MetaICL $74.9\%$ $\pm 0.9\%$ $74.2\%$ $\pm 0.6\%$ $83.0\%$ $\pm 0.1\%$ $77.8\%$ $\pm 0.8\%$ $82.2\%$ $\pm 0.3\%$
ABMLL-MetaICL (ours) $77.2\%$ $\pm 0.2\%$ $75.3\%$ $\pm 0.7\%$ $82.6\%$ $\pm 0.3\%$ $80.6\%$ $\pm 0.3\%$ $82.1\%$ $\pm 0.1\%$
Model Method cls-45 ECE $\downarrow$ cls-23 ECE $\downarrow$ NLI ECE $\downarrow$ Para ECE $\downarrow$ MCQA ECE $\downarrow$
Llama3 MetaICL $0.301$ $\pm 0.002$ $0.307$ $\pm 0.001$ $0.246$ $\pm 0.008$ $0.411$ $\pm 0.002$ $0.251$ $\pm 0.001$
Reptile-MetaICL $0.279$ $\pm 0.008$ $0.263$ $\pm 0.007$ $0.281$ $\pm 0.014$ $0.392$ $\pm 0.006$ $0.246$ $\pm 0.002$
ABMLL-MetaICL (ours) $0.241$ $\pm 0.000$ $0.271$ $\pm 0.003$ $0.207$ $\pm 0.006$ $0.402$ $\pm 0.017$ $0.244$ $\pm 0.006$
Qwen2 MetaICL $0.292$ $\pm 0.004$ $0.295$ $\pm 0.001$ $0.246$ $\pm 0.002$ $0.315$ $\pm 0.002$ $0.213$ $\pm 0.001$
Reptile-MetaICL $0.278$ $\pm 0.008$ $0.287$ $\pm 0.003$ $0.252$ $\pm 0.001$ $0.314$ $\pm 0.002$ $0.212$ $\pm 0.003$
ABMLL-MetaICL (ours) $0.249$ $\pm 0.005$ $0.284$ $\pm 0.002$ $0.227$ $\pm 0.015$ $0.288$ $\pm 0.012$ $0.217$ $\pm 0.001$
Table 3: Extension of Table 2 to LegalBench and ChemBench datasets, two domain applications that use ICL by default.
Model Method LegalBench Acc $\uparrow$ ChemBench Acc $\uparrow$ LegalBench ECE $\downarrow$ ChemBench ECE $\downarrow$
Llama3 MetaICL $77.1\%$ $\pm 0.1\%$ $55.0\%$ $\pm 0.4\%$ $0.265$ $\pm 0.002$ $0.403$ $\pm 0.001$
Reptile-MetaICL $79.1\%$ $\pm 0.2\%$ $54.6\%$ $\pm 0.4\%$ $0.261$ $\pm 0.004$ $0.393$ $\pm 0.003$
ABMLL-MetaICL (ours) $79.5\%$ $\pm 0.1\%$ $55.5\%$ $\pm 0.1\%$ $0.285$ $\pm 0.006$ $0.394$ $\pm 0.004$
Qwen2 MetaICL $75.7\%$ $\pm 0.1\%$ $53.4\%$ $\pm 0.0\%$ $0.291$ $\pm 0.004$ $0.417$ $\pm 0.001$
Reptile-MetaICL $78.8\%$ $\pm 0.1\%$ $58.1\%$ $\pm 0.1\%$ $0.252$ $\pm 0.003$ $0.394$ $\pm 0.001$
ABMLL-MetaICL (ours) $78.3\%$ $\pm 0.2\%$ $58.4\%$ $\pm 0.9\%$ $0.246$ $\pm 0.002$ $0.397$ $\pm 0.003$
5.2 Combining Meta-Learning with ICL
Meta-learning is a complementary approach to ICL, and being able to use meta-learning at scale makes it possible to combined these techniques. We compare using MetaICL (Min et al., 2022) and augmenting Reptile and ABMLL with in-context learning.
This section extends the evaluation datasets to two more domain applications with in-context learning - legal and chemistry domains (Guha et al., 2023; Mirza et al., 2024). For these evaluations, the training data is the same as the previous cls45 training split, a split that is generic for CrossFit and UnifiedQA. We hypothesize that even if the training data were distant from legal and chemistry domains, the weight adaptability encouraged by meta-learning should speed up few-shot learning for these domains.
5.2.1 Experimental Results - CrossFit, UnifiedQA, LegalBench, ChemBench
Meta-learning outperforms default LoRA to an even greater extent when combined with in-context examples, as demonstrated by the performance of ABMLL-MetaICL and Reptile-MetaICL (Table 2). ABMLL shows the strongest synergy with ICL, outperforming regular LoRA by over $10\%$ in accuracy on several datasets, and significantly outperforming regular MetaICL as well. The two meta-learning methods also generalize better to LegalBench and ChemBench (Table 3). These results suggests that meta-learning provides complementary gains to ICL, with the potential to improve performance in settings where ICL is used.
5.3 Model Pruning
Model pruning improves efficiency and probes robustness by removing redundant parameters, potentially eliminating spurious correlations (Ashkboos et al., 2024; Sun et al., 2023). Bayesian neural networks are known to be resource-efficient (Blundell et al., 2015). Inspired by this idea, we test the performance of ABMLL and benchmarks by setting a percentage of neurons in each layer embedding to zero, sorted by magnitudes of these neurons.
Table 5(b) shows that ABMLL is significantly more robust against this form of pruning than the other methods. This result suggests that ABMLL is robust and reliable, learning generalizable features that are not as tied to specific parameters. Although weight decay can improve pruning performance relative to no regularization, it hurts the unpruned setting and still underperforms ABMLL under pruning.
Method $0\%$ Pruned $1\%$ Pruned $10\%$ Pruned $20\%$ Pruned $30\%$ Pruned
(a) NLI.
Method $0\%$ Pruned $1\%$ Pruned $10\%$ Pruned $20\%$ Pruned $30\%$ Pruned
(b) Para.
5.4 Ablation Studies
Table 5 shows ABMLL performance across values of $\beta$ on cls-45. Results show that $\log_{10}\beta=0$, i.e., $\beta=1$, drops performance significantly. Thus, it is critical to balance reconstruction error and the KL terms that control how close the task-specific parameters $\phi_{i}$ are to global parameters $\theta$, and how close $\theta$ are to the prior $p(\theta)$. At the other extreme where the KL terms are too heavily tempered, the objective becomes degenerate, leading to poor performance ($\log_{10}\beta=-16$). In our experiments we searched for $\log_{10}\beta\in\{0,4,5,6,7,8,9,10\}$ with one random seed on cls-45 and identified the optimal value as $\log_{10}\beta=-8$. We use this setting for $\beta$ across all experiments.
In summary, while losses do not vary much with small changes in $\beta$, this study demonstrates that two extremes of $\beta$ drop performance significantly: either one where no balancing is used ($\beta=1$), or one that essentially drops the KL terms with an extremely small $\beta$.
Table 5: Validation accuracy and ECE across values of $\beta$ on cls-45. Lower $\beta$ means that the KL terms are more tempered, with pure maximization of data log likelihood in the limit.
$\log_{10}\beta$ 0 -5 -8 -16
Accuracy $\uparrow$ $64.5\%$ $70.4\%$ $75.2\%$ $60.6\%$
ECE $\downarrow$ $0.395$ $0.289$ $0.262$ $0.395$
6 Discussion
Our results show that a Bayesian approach to meta-learning for LoRA fine-tuning achieves strong generalization performance, measured in both accuracy and robustness, without a significant increase in computational cost. In the remainder of the paper we consider the implications of these results for broader questions about Bayesian methods and inductive biases for large language models. We also identify some of the limitations of our analyses and directions for future work.
Bridging the gap between Bayesian deep learning and LLMs.
Bayesian modeling offers a principled way to encode prior knowledge and represent uncertainty (Blei, 2014; Griffiths et al., 2008), and has been extensively studied in smaller neural networks (MacKay, 1995; Blundell et al., 2015). For LLMs, however, the computational cost of Bayesian methods has limited their use, despite their promise for uncertainty quantification, interpretability, and domain adaptation (Papamarkou et al., 2024). Our work helps bridge this gap by using LoRA to make Bayesian meta-learning practical for LLM fine-tuning.
Inductive bias in the era of LLMs.
In-context learning (Min et al., 2022; Chen et al., 2022) is a practical way to incorporate domain knowledge, but it can struggle when the relevant information is complex or unfamiliar (Dou et al., 2026). Gradient-based meta-learning of model weights is often more expressive, but substantially more expensive. Data synthesis offers another route to injecting inductive bias (McCoy et al., 2020). Our results suggest a scalable alternative: in-weights meta-learning through Bayesian LoRA adaptation. This approach improves generalization, can be combined with in-context learning, and maintains memory overhead that is constant in the number of tasks.
Meta-learning as a potential post-training stage for LLMs.
We have shown that meta-learning on the cls45 split provides strong performance on unseen tasks that are not closely related. An open problem to explore further is finding a stronger, larger-scale general-purpose dataset for metalearning in order to produce stronger few-shot performance on various unseen domain applications.
Conclusion.
Meta-learning is an effective method for supporting better generalization across datasets, but its demands on computation and memory can make it difficult to apply to large language models. We have shown how meta-learning can be applied to LLMs at scale by combining Amortized Bayesian Meta-Learning with Low-Rank Adaptation. This approach results in higher accuracy and across several benchmarks, and produces models that are more robust to pruning than existing methods for meta-learning in LLMs. Furthermore, being able to use meta-learning in these models allows us to demonstrate that it is complementary to in-context learning, allowing in-context examples to be used more effectively.
Appendix A Appendix
A.1 Algorithm
Full algorithm for ABMLL is given in Algorithm 1.
@@ A.1 Algorithm
  Run a step gradient descent to minimize w.r.t $\phi_{i}^{c}$$\mathbb{E}_{q_{\theta^{c}}(\phi_{i}^{c}|D_{i})}[\log p(D_{i}|\phi_{i}^{c})].$
@@
A.2 Additional Experiment Results